### v1

# Random local access for sampling $k$-SAT solutions

Abstract
We present a sublinear time algorithm that gives random local access to the uniform distribution over satisfying assignments to an arbitrary $k$-CNF formula $\Phi$, at exponential clause density. Our algorithm provides memory-less query access to variable assignments, such that the output variable assignments consistently emulate a single global satisfying assignment whose law is close to the uniform distribution over satisfying assignments to $\Phi$.
Such models were formally defined (for the more general task of locally sampling from exponentially sized sample spaces) in 2017 by Biswas, Rubinfeld, and Yodpinyanee, who studied the analogous problem for the uniform distribution over proper $q$-colorings. This model extends a long line of work over multiple decades that studies sublinear time algorithms for problems in theoretical computer science. Random local access and related models have been studied for a wide variety of natural Gibbs distributions and random graphical processes. Here, we establish feasiblity of random local access models for one of the most canonical such sample spaces, the set of satisfying assignments to a $k$-CNF formula.
1 Introduction
Efficiently sampling from a given exponentially sized solution space is a fundamental problem in computation. The quintessential such solution space in computer science is the collection of satisfying assignments to a given $k$-CNF (i.e., the solutions to a given SAT instance). Throughout, we let $\Phi$ denote an $n$-variable $(k,d)$-formula, which is a $k$-CNF formula such that each variable appears in at most $d$ clauses. We let $\Omega=\Omega_{\Phi}$ denote the set of satisfying assignments to $\Phi$, and let $\mu=\mu_{\Phi}$ denote the uniform distribution over $\Omega$. In this work, we view $k$ as a constant and take $n\to\infty$.
For $d\leq(2ek)^{-1}2^{k}$, it has long been known (e.g., by the Lovász local lemma introduced by Erdős and Lovász [6] in 1975) that any such $k$-CNF $\Phi$ has at least one satisfying assignment. Moreover, seminal work of Moser and Tardos [23] gave a simple algorithm that computes such a satisfying assignment in expected linear time. However, a similarly basic question is far more poorly understood: when is it possible to efficiently sample an (approximately uniformly) random satisfying assignment to $\Phi$?
The first breakthrough in this direction came from Moitra [21] who gave a deterministic, fixed-parameter tractable algorithm for approximately sampling from $\mu$ provided that $d\lesssim 2^{ck}$ for $c\approx 1/60$. This was followed up by work of Feng, Guo, Yin and Zhang [8], who used a Markov chain approach to give a polynomial time algorithm for sampling from $\mu$ provided that $d\lesssim 2^{k/20}$.
A natural question about such algorithms is whether one can adapt them to more efficiently answer local questions about a random satisfying assignment in sublinear time. For variable $v\in V$, let $\mu_{v}$ denote the marginal distribution on $v$ induced by $\mu$. When $n$ is large, one might wish to sample a small handful of variables $v$ from $\mu_{v}$ in $o(n)$ time, without computing an entire $\Omega(n)$-sized satisfying assignment $\sigma\sim\mu$. Traditionally, algorithms for counting the number of satisfying assignments to a given $k$-CNF proceed by computing marginal probabilities of variable assignments, a task that can be completed given local access to a random assignment. Therefore, sublinear time algorithms for answering local questions can also yield speedups in more general counting algorithms.
Ideally, such a random local access model should provide query access to variable assignments such that the output enjoys the following properties:
  1. (a)
    the model is consistent: queries for different variable assignments consistently emulate those of a single random satisfying assignment $\sigma\sim\mu$;
  2. (b)
    the model is sublinear: sampling variable assignments takes $o(n)$ time in expectation;
  3. (c)
    the model is memory-less: given the same initialization of randomness, answers to multiple, possibly independent queries for variable assignments are consistent with each other.
We give a more formal definition of random local access models in Section 2. Random local access models were formally introduced in recent work of Biswas, Rubinfeld, and Yodpinyanee [4], motivated by a long line of work over the past decades studying sublinear time algorithms for problems in theoretical computer science. The authors of [4] studied a different natural Gibbs distribution, the uniform distribution over proper $q$-colorings of a given $n$-vertex graph with maximum degree $\Delta$. By adapting previously studied classical and parallel sampling algorithms for $q$-colorings, they were able to construct a local access algorithm to a random proper $q$-coloring when $q\geq 9\Delta.$ The related problem of sampling partial information about huge random objects was pioneered much earlier in [13, 14]; further work in [24, 7] considers the implementation of different random graph models. Random local access and related (e.g., parallel computation) models have also been recently studied for several other random graphical processes and Gibbs distributions (cf. [4, 3, 22, 20]).
The above model for local access to random samples extends a rich line of work studying the more general local computation algorithms (LCAs), originally introduced by Rubinfield–Tamir–Vardi–Xie [25] and Alon–Rubinfield–Vardi–Xie [1]. A primary initial application of LCAs in [25] was locally constructing a satisfying assignment to a given $(k,d)$-formula $\Phi$ when $d\lesssim 2^{k/10}$. Similar to the desired properties of random local access, LCA implements query access to a large solution space in sublinear time using local probes. However, rather than sampling from a given distribution, LCA only provides local access to some consistent valid solution in the desired solution space that may otherwise be completely arbitrary in nature.
In this work, we present a sublinear time algorithm that gives random local access to the uniform distribution over satisfying assignments of an arbitrary $k$-CNF $\Phi$ at exponential clause density. The following is an informal version of our main result; a more formal version can be found at 4.7.
Theorem 1.1 (Main theorem: informal).
There exists a random local access algorithm $\mathcal{A}$, such that for any $(k,d)$-formula $\Phi=(V,\mathcal{C})$ that satisfies
$d\lesssim 2^{k/400}$ (1)
and any variable $v\in V$, $\mathcal{A}$ can be used to sample a single assignment $\widehat{\mu}_{v}\in\{0,1\}$ in $\mathrm{polylog}(n)$ time in expectation such that for all $c>0$,
$d_{\operatorname{\mathsf{TV}}}(\mathsf{Law}(\widehat{\mu}_{v}),\mu_{v})\leq\frac{1}{n^{c}}.$
Similar to [4], the proof of 1.1 adapts some of the algorithmic tools that came from the parallel/distributed sampling literature. The proof also builds upon the recent progress by Moitra [21] and Feng–Guo–Yin–Zhang [8] on sampling from bounded degree $k$-CNFs in polynomial time. The authors of [21, 8] both critically took advantage of a global variable marking (see 2.8) to better condition the marginal distributions of variables. Such an approach allows for a subset of the variable assignments to be sampled with ease; the resulting shattering of the solution space conditioned on such a partial assignment then allows one to efficiently complete the random satisfying assignment. These initial approaches have been extended and strengthened to nearly linear time algorithms that succeed for larger maximum variable degree in a flurry of recent work (cf. [21, 8, 18, 15, 11, 5, 10]).
More recently, the work of He, Wang and Yin [16] used a recursive sampling scheme to sample $k$-CNF solutions. Their work immediately gives a sublinear (in fact, expected constant time) algorithm for sampling the assignment of a single variable one time; however, this work does not immediately extend out of the box to a random local access model that enjoys the desired consistency and memory-less properties. Recursive sampling schemes have also been recently used to make progress on analyzing and designing fast sampling algorithms for a variety of Gibbs distributions (cf. [16, 5, 2]).
An immediate roadblock to adapting global sampling strategies to random local access to $k$-CNF solutions is that the vast majority of the aforementioned existing algorithms critically use some global information – like a variable marking, particular ordering of variables, or other global state compression – as an input to make algorithmic decisions, necessitating a substantive departure from these approaches for any hope of local access adaptation. In this work, we overcome these obstacles by carefully selecting and sampling a small number of auxiliary variable assignments on an as-needed basis, and showing that bad cases can be reduced to small brute force calculations after a sublinear number of steps. Thus our proof of 1.1 requires a novel adaptation of sampling techniques that require global information or complete passes of the variable set to the local setting, showing that we can perform analogous operations locally or circumvent them altogether in the course of locally sampling a variable assignment.
Acknowledgements
We would like to thank Kuikui Liu and Ronitt Rubinfeld for helpful discussions, suggestions, and feedback on an early draft. NM was supported by the Hertz Graduate Fellowship and the NSF Graduate Research Fellowship Program #2141064.
2 Preliminaries
Notation
Throughout, we let $\Phi=(V,\mathcal{C})$ be a $k$-CNF on variable set $V=\{v_{1},\ldots,v_{n}\}$ with associated collection of clauses $\mathcal{C}=\{C_{1},\ldots,C_{m}\}$. We further assume that $\Phi$ is a $(k,d)$-formula; that is, each variable $v_{i}$ appears in at most $d$ clauses. For every clause $C\in\mathcal{C}$, let $\operatorname{vbl}(C)$ denote the collection of variables in the clause $C$. We further define $\Delta:=\max_{C\in\mathcal{C}}|\{C^{\prime}\in\mathcal{C}:\operatorname{vbl}(C)\cap\operatorname{vbl}(C^{\prime})\neq\emptyset\}|$, so in particular $\Delta\leq kd$.
In the regime we work in, we assume $k$ is a large fixed constant and view $n\to\infty.$ We use $f\lesssim g$ to denote that there is some constant $C$ (not depending on $k$) such that $f\leq Cg$. We also use the standard asymptotic notation ($O$, $o$, $\Omega$, $\omega$, $\Theta$), where when not specified, we assume these are in the $n\to\infty$ limit and in these notations, we ignore any dependencies on $k$. We use $\mathsf{Law}(X)$ to denote the underlying distribution of a random variable $X$.
We let $\Omega=\Omega_{\Phi}\subseteq\{0,1\}^{n}$ denote the collection of satisfying assignments to $\Phi$ and let $\mu=\mu_{\Phi}$ denote the uniform distribution over $\Omega$. Throughout, we suppress the $\Phi$ subscripts when the formula is clear from context. We now introduce a few more concepts associated to a $k$-CNF formula that will be used later.
Definition 2.1 (Dependency graph).
Given a $(k,d)$-formula $\Phi$, we define the following auxiliary (hyper)graphs.
  1. 1.
    Let $H_{\Phi}=(V,\mathcal{E})$ be the dependency hypergraph (with multiple edges allowed), where $V$ is the set of variables and =$\mathcal{E}=\{\operatorname{vbl}(C):C\in\mathcal{C}(\Phi)\}$ is the collection of clauses of $\Phi$ viewed as $k$-sets of variables;
  2. 2.
    $G_{\Phi}$ be the clause dependency graph of $\Phi$, where $V(G_{\Phi}):=\mathcal{C}$ is the set of clauses of $\Phi$, and two clauses $C,C^{\prime}\in\mathcal{C}$ are adjacent in $G_{\Phi}$ if they share a variable.
Definition 2.2 (Partial assignment).
Given $(k,d)$-formula $\Phi=(V,\mathcal{C})$, let
$\mathcal{Q}^{*}:=\{0,1,\star,\circ\}^{V}$
denote the space of partial assignments with the following symbology. Given a partial assignment $\sigma\in\mathcal{Q}^{*}$, each variable $v\in V$ is classified as follows:
  • $\sigma(v)\in\{0,1\}$ means that $v$ is accessed by the algorithm and assigned with $\sigma(v)\in\{0,1\}$;
  • $\sigma(v)=\star$ means that $v$ is just accessed by the algorithm but unassigned yet with any value;
  • $\sigma(v)=\circ$ means that $v$ is unaccessed by the algorithm and hence unassigned with any value.
We sometimes use $\emptyset$ to denote the empty assignment (i.e., $\sigma(v)=\circ$ for all $v\in V$). For every $\sigma\in\mathcal{Q}^{*}$, we use $\Lambda(\sigma)\subseteq V$ to denote the set of variables that are assigned and accessed in $\sigma$, namely
$\Lambda(\sigma):=\{v\in V:\sigma(v)\in\{0,1\}\}.$
Given $\sigma\in\mathcal{Q}^{*}$, variable $v\in V$ and $x\in\{0,1,\star,\circ\}$, we further denote by $\sigma_{v\leftarrow x}$ the partial assignment obtained from $\sigma$ by replacing $\sigma(v)$ with $x$. We say that $\sigma$ is feasible if it can be extended to a satisfying assignment to $\Phi$.
Definition 2.3 (Reduced formula on partial assignment).
Let $\Phi$ be a $(k,d)$-formula. Given a partial assignment $\sigma$ on $\Lambda:=\Lambda(\sigma)$, let $\Phi^{\sigma}$ be the result of simplifying $\Phi$ under $\sigma$. That is, define $\Phi^{\sigma}:=(V^{\sigma},\mathcal{C}^{\sigma})$ where
  • $V^{\sigma}=V\setminus\Lambda$,
  • $\mathcal{C}^{\sigma}$ is obtained from $\mathcal{C}$ by removing all clauses that have been satisfied under $\sigma$ and removing any appearance of variables in $\Lambda$.
Let $H_{\Phi}^{\sigma}$ be the associated (not necessarily $k$-uniform) hypergraph to $\Phi^{\sigma}$, and let $G_{\Phi}^{\sigma}$ be the associated clause dependency graph to $\Phi^{\sigma}$. For variable $v\in V\setminus\Lambda$, let $\Phi_{v}^{\sigma}$ denote the connected component of $\Phi^{\sigma}$ to which $v$ belongs.
Definition 2.4 (Marginal distribution).
For an arbitrary set of variables $S\subseteq V$, let $\mu_{S}$ be the marginal distribution on $S$ induced by $\mu$, so that
$\mu_{S}(\tau)=\sum_{\tau^{\prime}\in\{0,1\}^{n}:\tau^{\prime}|_{S}=\tau}\mu(\tau^{\prime})\qquad\forall\tau\in\{0,1\}^{S}.$
When $S=\{v\}$ is a single vertex, we let $\mu_{v}=\mu_{\{v\}}$. Furthermore, given some partial assignment $\sigma\in\{0,1\}^{\Lambda}$ for $\Lambda\subseteq V$, if $S\cap\Lambda=\emptyset$, we let $\mu_{S}^{\sigma}(\cdot):=\mu_{S}(\cdot\mid\sigma)$ be the marginal distribution on $S$ conditioned on the partial assignment $\sigma$. When $S=V\setminus\Lambda$, we simply use $\mu^{\sigma}$ to denote $\mu^{\sigma}_{V\setminus\Lambda}$.
2.1 The random local access model and local computation algorithms
One of the most widely studied models of local computation are local computation algorithms (LCAs) introduced in [1, 25]. Given a computation problem $F$, an LCA (in an online fashion) provides the $i$-th bit to some consistent solution $F$ in sublinear time. As originally defined, local computation algorithms need not be memory-less; in other words, the output solution can depend on the order of queried bits. However, several follow-up works have given query-oblivious or memory-less analogues of LCAs for a variety of natural problems. Such models are the non-random-sampling version of the random local access models we study here.
In this work, we construct a query-oblivious LCA for an intermediate step in our construction of a random local access model (described in more detail in Section 2.3). We thus precisely define both LCAs and random local access models below.
Definition 2.5 (Local computation algorithm).
Given an object family $(\Pi,\Omega)$ with input $\Pi$, sample space $\Omega\subseteq\Sigma^{n}$ (for some alphabet $\Sigma$), a $(t(n),\delta(n))$-local computation algorithm (LCA) provides an oracle $\mathcal{A}$ that implements query access to some arbitrary $\sigma\in\Omega$ that satisfies the following conditions:
  • $\mathcal{A}$ has query access to the input description $\Pi$ and to a tape of public random bits $\mathbf{R}$.
  • $\mathcal{A}$ gets a sequence of queries $i_{1},\ldots,i_{q}$ for any $q>0$, and after each query $i_{j}$, it produces an output $\sigma_{i_{j}}$ such that the outputs $\sigma_{i_{1}},\ldots,\sigma_{i_{q}}$ are consistent with some $\sigma\in\Omega$.
  • The probability of success over all $q$ queries is at least $1-\delta(n)$ (where $\delta(n)<1/3$).
  • The expected running time of $\mathcal{A}$ on any query is at most $t(n)$, which is sublinear in $n$.
We say that $\mathcal{A}$ is a $(t(n),\delta(n))$-strongly local computation algorithm if $t(n)\lesssim\log^{c}(n)$ for some constant $c>0.$ We further say that $\mathcal{A}$ is query-oblivious if the outputs of $\mathcal{A}$ do not depend on the order of the queries but depend only on $\Pi$ and $\mathbf{R}$.
Motivated by the above, we give a formal definition of the random local access model introduced in [4].
Definition 2.6 (Random local access).
Given a random object family $(\Pi,\Omega,\mu)$ with input $\Pi$, sample space $\Omega\subseteq\Sigma^{n}$ (with alphabet $\Sigma$), and distribution $\mu$, a $(t(n),\delta(n))$-random local access implementation of a family of query functions $\{F_{1},F_{2},\ldots\}$ provides an oracle $\mathcal{A}$ with the following properties:
  • $\mathcal{A}$ has query access to the input description $\Pi$ and to a tape of public random bits $\mathbf{R}.$
  • For a given input $\Pi$, upon being queried with $F_{i}$, the oracle with probability $1-o(1)$ uses at most $t(n)$ resources (where $t(n)$ is sublinear in $n$) to return the value $\mathcal{A}(\Pi,\mathbf{R},F_{i}(Y))$ for some specific $Y\in\Omega$.
  • The choice of $Y\in\Omega$ only depends on $\Pi$ and $\mathbf{R}$.
  • The distribution of $Y$ over the randomness $\mathbf{R}$ satisfies
    $\|\mathsf{Law}(Y)-\mu\|_{\operatorname{\mathsf{TV}}}=\frac{1}{2}\sum_{\omega\in\Omega}\|\mathbb{P}(Y=\omega)-\mu(\omega)\|<\delta(n),$
    where $\delta(n)\lesssim\frac{1}{n^{c}}$ for constant $c>1$.
In other words, if $\mathcal{A}$ is a random local access oracle for a set of queries, then when provided the same input $\Pi$ and the same random bits $\mathbf{R}$, it must provide outputs that are consistent with a single choice of $Y$ regardless of the order and content of these queries.
Remark 2.7.
In this work, we do not study or seek to optimize the memory usage of our algorithms. However, there is also a rich literature studying space-efficient and parallelizable local models (e.g., [1, 12, 19]).
2.2 Marking and shattering
As noted in the introduction, the Lovász local lemma guarantees the existence of a satisfying assignment to any $(k,d)$-formula $\Phi$ if $d\leq(2ek)^{-1}2^{k}.$ Furthermore, the Moser-Tardos algorithm [23] gives a simple linear-time algorithm to construct such a satisfying assignment in the above $(k,d)$-regime:
[…]
  • While there exists a clause $C\in\mathcal{C}(\Phi)$ that is violated by the current assignment, resample $\operatorname{vbl}(C)$ uniformly at random.
The Lovász local lemma can be made quantitative, showing that not only is there some satisfying assignment to $\Phi$ if $d\leq(2ek)^{-1}2^{k}$, but both that there are exponentially many such satisfying assignments and that the marginal distributions $\mu_{v}$ are approximately uniform with $\mu_{v}(1)\leq\frac{1}{2}\exp(1/k)$ (see [21, 8]). Such local uniformity is critical to the success of algorithms that approximately sample from $\mu$. In their breakthrough work, Moitra [21] noted that this local uniformity continues to hold for conditional distributions $\mu^{\sigma}$ provided that each clause has a sufficiently large number of free variables under partial assignment $\sigma$. This motivates the idea of a marking, as introduced in [21], which is a careful selection (via the Lovász local lemma) of a linear sized subset of variables $\mathcal{M}\subseteq V$ that has the following properties:
  • For every clause $C$, $\mathcal{M}\cap\operatorname{vbl}(C)\gtrsim k$. Having a large number of marked variables in each clause would hopefully result in a desired shattering condition, namely, we can sample a partial assignment $\sigma:\mathcal{M}\to\{0,1\}$ on this marking that partitions the original formula into sufficiently small connected components.
  • For every clause $C$, $\operatorname{vbl}(C)\backslash\mathcal{M}\gtrsim k$. This large non-intersection preserves the local uniformity of the marginal distributions of the variables in $\mathcal{M}$.
The general strategy of arguments following the marking approach is to show that it is “easy” to sample a partial assignment $\sigma:\mathcal{M}\to\{0,1\}$, and moreover, conditioned on any such assignment, the dependency graph $G_{\Phi^{\sigma}}$ is very likely to shatter into sufficiently small connected components such that the remaining variable assignments can be efficiently sampled from the conditional distribution by solving a smaller instance of the original problem.
We make these notions precise and give a simple condition under which there exists a linear-time algorithm to construct such a marking.
Definition 2.8 ($\alpha$-marking).
Given $(k,d)$-formula $\Phi$ and constant $\alpha\in(0,1/2)$, we say that $\mathcal{M}\subseteq V$ is an $\alpha$-marking if for every $C\in\mathcal{C}$, $|\operatorname{vbl}(C)\cap\mathcal{M}|\geq\alpha k$ and $|\operatorname{vbl}(C)\backslash\mathcal{M}|\geq\alpha k$.
Lemma 2.9 ([8]).
Given $(k,d)$-formula $\Phi$ and constant $\alpha\in(0,1/2)$, if $d\lesssim\frac{1}{\mathrm{poly}(k)}2^{k\frac{(1-2\alpha)^{2}}{6\ln 2}}$, then there exists an $\alpha$-marking, and one such marking can be found via the Moser-Tardos algorithm.
We further note that the desired shattering occurs with high probability conditioned on a partial assignment on the marked variables that follows the marginal distribution $\mu_{\mathcal{M}}$.
Definition 2.10 (Shattering).
Given $(k,d)$-formula $\Phi$. We say that assignment $\sigma:\mathcal{M}\to\{0,1\}$ $\delta$-shatters $\Phi$ if for all $v\in V$, the connected component in $G_{\Phi}^{\sigma}$ that contains $v$ has size $\lesssim dk\log(n/\delta).$
Lemma 2.11 ([8]).
If $\Phi$ is a $(k,d)$-formula and $\mathcal{M}$ is a valid $\alpha$-marking for constant $\alpha\in(0,1/2)$, then for $2^{\alpha k}\gtrsim d^{3}k^{3}$, we have
$\mathbb{P}_{\tau\overset{\mathrm{\mathbf{R}}}{\leftarrow}\mu_{\mathcal{M}}}[\tau\,\delta\text{-shatters }\Phi]\geq 1-\frac{\delta}{n^{2}}.$
In this work, we locally, greedily construct an $\alpha$-marking $\mathcal{M}$ using a local implementation of Moser-Tardos. We will further argue that because of the shattering property, we can locally compute the connected component of $G_{\Phi}^{\sigma}$ for some $\sigma\sim\mu_{\mathcal{M}}$ that contains a given vertex, without having to actually assign linearly many vertices.
2.3 Query-oblivious LCA for marking
Per above, we construct a query-oblivious LCA, $\mathsf{IsMarked}(\cdot)$, where for a $(k,d)$-formula $\Phi$ and a given range of $\alpha\in(0,1)$, $\mathsf{IsMarked}(\cdot)$ can take in any variable $v\in V$ and output either 0 or 1 indicating whether $v$ is in some $\alpha$-marking of $V$. Moreover, $\mathsf{IsMarked}(\cdot)$ takes sublinear time and when queried on all of $V$, gives a consistent $\alpha$-marking of $V$.
Theorem 2.12.
Let $\Phi=(V,\mathcal{C})$ be a $(k,d)$-formula. Suppose there exist constants $1/2<\beta_{1}<\beta_{2}<1-\alpha$ that satisfy the following conditions:
$4\alpha<2(1-\beta_{2})<1-\beta_{1},$
$16k^{4}d^{5}\leq 2^{(\beta_{1}-h(1-\beta_{1}))k},$
$16k^{4}d^{5}\leq 2^{(\beta_{2}-\beta_{1})k-h(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}})(1-\beta_{1})k},$ (2)
$\delta\mapsto(\beta_{2}-\delta)-h(\frac{\beta_{2}-\delta}{1-\delta})(1-\delta)\text{ is decreasing on $0\leq\delta\leq\beta_{1}$,}$
$2e(kd+1)<2^{(1-h(\frac{\alpha}{1-\beta_{2}}))(1-\beta_{2})k}.$
(Here $h(\cdot)$ denotes the binary entropy function $x\mapsto-x\log_{2}(x)-(1-x)\log_{2}(1-x)$.)
Then there exists a $\mathrm{polylog}(n)$ time local computation algorithm $\mathsf{IsMarked}(\cdot)$ which, given any variable $v\in V$, returns an assignment $s_{v}\in\{0,1\}$ denoting whether $v$ is contained in an $\alpha$-marking of $\Phi$. Moreover, with probability at least $1-1/n$, the response for every $v\in V$ gives a consistent $\alpha$-marking of $\Phi$.
The construction of $\mathsf{IsMarked}(\cdot)$ and the verification that it is a query-oblivious LCA is inspired by the approach used in [1] to obtain an oblivious LCA for hypergraph $2$-coloring. Thus, we defer the proof of Theorem 2.12 to Appendix A.
3 A random local access algorithm for $k$-CNF solutions
In this section, we introduce an algorithm (Algorithm 1) that locally accesses the assignment of a variable $v\in V$ in a uniformly random solution $\sigma$ to $\Phi$. Overall, given $k$-CNF formula $\Phi=(V,\mathcal{C})$ and variable $v\in V$, the algorithm performs the following:
  1. 1.
    Locally decides whether $v$ lies an $\alpha$-marking of $\Phi$ ($\mathsf{IsMarked}(v)$), such that the response for every $v\in V$ gives a consistent $\alpha$-marking $\mathcal{M}\subseteq V$.
  2. 2.
    If $v$ is marked, then we sample $\sigma(v)$ by computing $\mathsf{MarginSample}(\Phi,\emptyset,v)$ (adapted from [16]), which may make further recursive calls to sample other marked variables.
  3. 3.
    If $v$ is not marked, then we perform a depth-first search starting from $v$ in the reduced dependency hypergraph to compute the restriction of $\sigma$ on the marked variables. We start from $\sigma=\emptyset$; for every marked variable $w$ we encounter that we have not yet sampled, we compute $\mathsf{MarginSample}(\Phi,\sigma,w)$ and update $\sigma(w)$, to eventually obtain a connected component $\Phi_{v}^{\sigma}$ that contains $v$ (that we will prove is polylogarithmic in size with high probability). We then perform a $\mathsf{UniformSample}$ algorithm (Theorem 4.3) to sample a uniformly random satisfying assignment $\sigma$ for this connected component, and output $\sigma(v)$.
Algorithm 1 is our main routine; it has several subroutines which we describe in detail below.
Throughout the algorithm, we use public randomness $\mathbf{R}$ to fix a random global ordering of the variables $v_{1},\dots,v_{n}$ (e.g., by assigning an independent random value in $[0,1]$ to each $v_{i}$), such that one can locally query the relative ordering between any two variables in constant time.
Algorithm 1 The sampling algorithm
Input: $k$-CNF formula $\Phi=(V,\mathcal{C})$ and variable $v\in V$
      Output: random assignment $\sigma(v)\in\{0,1\}$
1:Compute $\mathsf{IsMarked}(v)$
2:if $\mathsf{IsMarked}(\Phi,v)$ then
3:     return $\mathsf{MarginSample}(\Phi,\emptyset,v)$
4:else
5:     $\Phi_{v}\leftarrow\mathsf{Conn}(\Phi,\emptyset,v)$
6:     $\sigma\leftarrow\mathsf{UniformSample}(\Phi_{v})$
7:     return $\sigma(v)$
We define (as in [16])
$\theta:=\frac{1}{2}-((1-2e\xi)^{-dk}-1)-(32ek^{2}d)^{-1},\quad\xi=\frac{1}{72e^{2}k^{3}d^{2}}.$ (3)
For a sufficiently sparse $(k,d)$-CNF $\Phi$, $\theta$ will be a pointwise lower bound on the marginal probabilities $\mu_{v}(\cdot)$. The basis of our local sampler will be the observation that $\theta\approx\frac{1}{2}$. Thus, we can sample the assignment of $v$ from $\mu_{v}$ via the following procedure: with probability $2\theta$, we uniformly at random select $v$ from $\{0,1\}$, and with probability $1-2\theta$, we sample from the overflow distribution $\mathcal{D}:=\frac{\mu_{v}^{\sigma}-\theta}{1-2\theta}$. Note that the first step can be completed in $O(1)$ time without attempting to sample any other variable spin. Thus, since $2\theta$ is relatively close to $1$, we can hope to locally sample $v$ without sampling too many other variable assignments. We make this idea more precise below with the $\mathsf{MarginSample}$ (Algorithm 2) and $\mathsf{MarginOverflow}$ (Algorithm 3) algorithms, adapted from [16].
Algorithm 2 $\mathsf{MarginSample}(\Phi,\sigma,v)$
Input: $k$-CNF formula $\Phi=(V,\mathcal{C})$, $v\in V$ with $\mathsf{IsMarked}(v)=1$, partial assignment $\sigma$
      Output: random assignment $\sigma(v)\in\{0,1\}$
[…]
2:if $r_{v}<2\theta$ then
[…]
4:else
5:     $\sigma(v)\leftarrow\mathsf{MarginOverflow}(\Phi,\sigma_{v\leftarrow\star},v)$
6:return $\sigma(v)$
Algorithm 3 $\mathsf{MarginOverflow}(\Phi,\sigma,v)$
Input: $k$-CNF formula $\Phi=(V,\mathcal{C})$, $v\in V$ with $\mathsf{IsMarked}(v)=1$, partial assignment $\sigma$
      Output: random assignment $\sigma(v)\in\{0,1\}$
1:Draw $u\leftarrow\mathsf{NextVar}(\sigma)$ (3.1)
2:if $u\neq\perp$ then
[…]
4:     if $r_{u}<2\theta$ then
[…]
6:     else
7:         $\sigma(u)\leftarrow\mathsf{MarginOverflow}(\Phi,\sigma_{u\leftarrow\star},u)$      $\sigma(v)\leftarrow\mathsf{MarginOverflow}(\Phi,\sigma,v)$
8:else
9:     $\Phi^{\sigma}_{v}\leftarrow\mathsf{Conn}(\Phi,\sigma,v)$
10:     $\sigma(v)\leftarrow\mathsf{UniformSample}(\Phi^{\sigma}_{v})$
11:return $\sigma(v)$
A major difference of our algorithm from that of [16] is that rather than going through all variables in order, we restrict calls to $\mathsf{MarginSample}$ and $\mathsf{MarginOverflow}$ to marked variables and attempt to only sample variables in a single marked component, which results in a different analysis of the associated sampling algorithm.
Definition 3.1 ($\mathsf{NextVar}$).
Given a $(k,d)$-formula $\Phi=(V,\mathcal{C})$ and partial assignment $\sigma$, let $V_{\star}\subseteq V\backslash\Lambda(\sigma)$ be the collection of variables with $\sigma(v)=\star$. Recall that =$\mathcal{M}=\{v\in V:\mathsf{IsMarked}(v)=1\}$. Define
$\partial V_{\star}=\{u\in\mathcal{M}\backslash(\Lambda(\sigma)\cup V_{\star})\mid\text{ there exist }C\in\mathcal{C},v\in V_{\star}\text{ such that }u,v\in\operatorname{vbl}(C)\}.$
Then, given a total ordering of $V$ (which we assume we have local access to by considering a random ordering defined by the public random bits $\mathbf{R}$), we let
$\mathsf{NextVar}(\sigma)=\begin{cases}v_{i}\in\partial V_{\star}\text{ with smallest $i$}&V_{\star}\neq\emptyset\\ \perp&\text{else}\end{cases}.$
Remark 3.2.
In words, $\mathsf{NextVar}(\sigma)$ is the smallest unaccessed marked variable (in the fixed global ordering of $V$ determined by $\mathbf{R}$) that shares a clause with some accessed but unassigned variable in the partial assignment $\sigma$. In the algorithm, $\mathsf{NextVar}(\sigma)$ can be found by querying all unaccessed marked neighbors of $V_{\star}$ and identifying the smallest element via local comparisons.
Algorithm 4 $\mathsf{Conn}(\Phi,\sigma,v)$
Input: CNF formula $\Phi=(V,\mathcal{C})$, $v\in V$, partial assignment $\sigma$
      Output: random assignment $\sigma(v)\in\{0,1\}$
1:Initialize $\mathcal{S}=\{C^{\sigma}\in\mathcal{C}^{\sigma}\mid v\in\text{vbl}(C^{\sigma})\}$ as a stack
2:while $|\mathcal{S}|>0$ do
3:     Pop $C^{\sigma}$ off of $\mathcal{S}$
4:     for $w\in\operatorname{vbl}(C^{\sigma})$ do
5:         if $\mathsf{IsMarked}(w)$ then
6:              $\sigma(w)\leftarrow\mathsf{MarginSample}(\Phi,\sigma,w)$
7:              Remove from $\mathcal{S}$ any clauses satisfied by $\sigma$
8:         if $C^{\sigma}=1$ then
9:              break
10:     if $C^{\sigma}\neq 1$ then
11:         for $w\in\operatorname{vbl}(C^{\sigma})$ do
12:              Push $\{(C^{\prime})^{\sigma}\mid C^{\prime}\in\mathcal{C},w\in\operatorname{vbl}((C^{\prime})^{\sigma}),(C^{\prime})^{\sigma}\neq 1\}$ onto $\mathcal{S}$               return $\Phi_{v}^{\sigma}$
Algorithm 1 contains several subalgorithms that are formally given above. We describe more colloquially how these algorithms work and are interconnected below. To sample the assignment of some $v\in V$, we first use $\mathsf{IsMarked}(\cdot)$ as in Theorem 2.12 to determine whether $v$ lies in a given $\alpha$-marking $\mathcal{M}$. The next steps of our sampling depend on this boolean outcome.
At the beginning of Algorithm 1, we have a target variable $v$ we wish to sample and initialize partial assignment $\sigma=\emptyset$. We first compute $\mathsf{IsMarked}(v)$. If $\mathsf{IsMarked}(v)=1$, then we perform $\mathsf{MarginSample}(\Phi,\sigma,v)$ (Algorithm 2) to sample the assignment of $v$. With probability $2\theta$ we uniformly at random select $v$ from $\{0,1\}$, and with probability $1-2\theta$ we sample from the overflow distribution $\mathcal{D}:=\frac{\mu_{v}^{\sigma}-\theta}{1-2\theta}$ by calling $\mathsf{MarginOverflow}(\Phi,\sigma_{v\leftarrow\star},v)$ (Algorithm 3). The notation $v\leftarrow\star$ means that we will temporarily mark $v$ as an accessed by unassigned variable (2.2) and sample its value in the end.
Going into $\mathsf{MarginOverflow}(\Phi,\sigma_{v\leftarrow\star},v)$, we will use $\mathsf{NextVar}(\cdot)$ to determine the smallest unaccessed marked neighbor of $v$ (call it $w$). Again, with probability $2\theta$ we randomly select $w$ from $\{0,1\}$, and with probability $1-2\theta$, we sample $w$ from the overflow distribution $\mathsf{MarginOverflow}(\Phi,\sigma_{w\leftarrow\star},u)$. We will show in Proposition 4.4 that, with high probability, this process terminates after a polylogarithmic number of $\mathsf{MarginOverflow}$ iterations. After that we compute $\Phi^{\sigma}_{v}$, the connected component that contains $v$ under the current partial assignment $\sigma$. Finally we sample from $\Phi^{\sigma}_{v}$ using a uniform sampling algorithm (Theorem 4.3) and output its $v$ coordiante, concluding the procedure $\mathsf{MarginSample}(\Phi,\emptyset,v)$.
Otherwise, if $v$ is not marked in $\mathcal{M}$, then we compute $\mathsf{Conn}(\Phi,\emptyset,v)$ (Algorithm 4), the connected component in $\Phi^{\sigma}$ that contains $v$ for $\sigma\sim\mu_{\mathcal{M}}$ some random assignment that we locally construct using depth first search starting from $v$. We initialize $\sigma=\emptyset$ and update $\sigma$ along the way by sampling any marked variables we encounter. Thus, we are able to locally construct the relevant portion of a desired, consistent partial assignment. Shattering results (Lemma 2.11) will give us a polylogarithmic-sized component in $\mathsf{Conn}$ and thus polylogarithmic running time of both $\mathsf{Conn}$ and the subsequent call to $\mathsf{UniformSample}$ (Theorem 4.3). We make this analysis more precise in subsequent sections.
4 Analysis
We verify the correctness and efficiency of Algorithm 1 by analyzing each of its steps. The $\mathsf{IsMarked}$ subroutine has been discussed in Section 2 (Theorem 2.12). In this section, we finish the analysis of $\mathsf{MarginSample}$, $\mathsf{Conn}$ and $\mathsf{UniformSample}$.
4.1 Correctness
Conditioned on $\mathsf{IsMarked}$ returning a consistent marking $\mathcal{M}$ over all $v\in V$ (which by Theorem 2.12 occurs with probability $1-1/n$), we show that with probability 1, $(\mathsf{MarginSample}(v))_{v\in\mathcal{M}}$ follows the marginal distribution $\mu_{\mathcal{M}}$ and computes a variable sample in expected $\mathrm{polylog}(n)$ time.
Correctness of the $\mathsf{MarginSample}$ and $\mathsf{MarginOverflow}$ routines will extend by an identical argument to the corresponding algorithms in [16], since we condition on a consistent marking and compute the same connected component but locally. More precisely, we apply the following results.
Theorem 4.1 ([16, Theorem 5.5]).
We have the following pair of correctness results:
  1. (1)
    Suppose that $(\Phi,\sigma,v)$ satisfies the following trio of conditions:
    • $\Phi=(V,\mathcal{C})$ is a $(k,d)$-formula where $v\in V$ and $\sigma$ is a feasible partial assignment;
    • For every clause $C\in\mathcal{C}$,
      $\mathbb{P}_{\mu}[\text{$C$ violated}\mid\sigma]\leq 2\xi,$
      for $\xi$ defined in 3;
    • The variable $v$ is not part of partial assignment $\sigma$ (i.e., $v\notin\Lambda(\sigma)$).
    Then, $\mathsf{MarginSample}(\Phi,\sigma,v)$ terminates with probability $1$, and it returns a random value $x\in\{0,1\}$ distributed as $\mu_{v}^{\sigma}$ upon termination.
  2. (2)
    Suppose that $(\Phi,\sigma,v)$ satisfies the following trio of conditions:
    • $\Phi=(V,\mathcal{C})$ is a $(k,d)$-formula where $v\in V$ and $\sigma$ is a feasible partial assignment;
    • For every clause $C\in\mathcal{C}$,
      $\mathbb{P}_{\mu}[\text{$C$ violated}\mid\sigma]\leq 2\xi,$
      for $\xi$ defined in 3;
    • The variable $v$ is not part of partial assignment $\sigma$ and furthermore has been given assignment $v\leftarrow\star$.
    Then, $\mathsf{MarginOverflow}(\Phi,\sigma,v)$ terminates with probability $1$, and it returns a random value $x\in\{0,1\}$ distributed as the overflow distribution $\mathcal{D}:=\frac{\mu_{v}^{\sigma}-\theta}{1-2\theta}$ upon termination.
With Theorem 4.1, we can verify the correctness and consistency of $\mathsf{MarginSample}(\Phi,\emptyset,v)$ for each $v\in\mathcal{M}$.
Proposition 4.2.
Let $\Phi$ be a $(k,d)$-formula and $\alpha\in(0,1)$ be a constant such that $d,\alpha$ satisfy the conditions of Theorem 2.12, and
$2^{4}\cdot d^{5}\cdot k^{6}\cdot 2^{-\alpha k}\leq\frac{1}{150e^{3}}.$ (4)
Given a public string of random bits $\mathbf{R}$, suppose that $(\mathsf{IsMarked}(\Phi,v))_{v\in V}$ is an oblivious LCA and gives a valid $\alpha$-marking $\mathcal{M}\subseteq V$. Then, $(\mathsf{MarginSample}(\Phi,\emptyset,v))_{v\in\mathcal{M}}$ is memory-less and follows the distribution $\mu_{\mathcal{M}}$.
Proof.
Consider some public randomness $\mathbf{R}$. This determines a global ordering of variables and induces a marking given by $(\mathsf{IsMarked}(v))_{v\in V}$. We condition on the outcome that $(\mathsf{IsMarked}(v))_{v\in V}$ gives a valid $\alpha$-marking $\mathcal{M}\subseteq V$, where =$\mathcal{M}=\{v_{1},\ldots,v_{s}\}$ with $v_{1}<\cdots<v_{s}$ with respect to the random global ordering.
Consider any marked variable $v\in\mathcal{M}$. To sample the assignment of $v$, $\mathsf{MarginSample}(\Phi,\emptyset,v)$ and subsequent recursive calls to $\mathsf{MarginOverflow}$ only attempt to sample the assignments of marked variables encountered along the algorithm. Thus, for any clause $C\in\mathcal{C}$, we have at least $\alpha k$ unmarked variables and thus
$\mathbb{P}_{\mu}[\text{$C$ violated}\mid\sigma]\leq 2^{-\alpha k}\leq 2\xi,$
which holds by our choice of $\xi$ in 3 and $d,\alpha$ in 4. This guarantees that in every call to $\mathsf{MarginSample}$ and $\mathsf{MarginOverflow}$, the conditions for both properties in Theorem 4.1 are satisfied.
We now show consistency of the algorithm, i.e, the outputs $(\mathsf{MarginSample}(\Phi,\emptyset,v))_{v\in\mathcal{M}}$ jointly follow the distribution $\mu_{\mathcal{M}}$. Whenever we compute the assignment of some $v\in\mathcal{M}$, we recompute the assignments of all variables $w$ whose assignment $v$ depends on. Computing the assignment of $w$ involves two steps: (a) first checking where $r_{w}$ falls relative to $\theta_{w}$, (b) calling $\mathsf{MarginOverflow}$ if we cannot immediately assign $w$. Step (a) only relies on the bits of randomness corresponding to $r_{w}$ from the shared randomness $\mathbf{R}$ and is independent of the order in which the variables are processed. Regarding (b), the global ordering and consistency of the marking (Theorem 2.12) ensure that variables are processed in a consistent order and thus retains the desired obliviousness property of the local sampler. ∎
Given a feasible partial assignment $\sigma$ to a satisfiable $(k,d)$-formula $\Phi$, $\mu_{v}^{\sigma}$ is independent of any variable that does not appear in $\Phi_{v}^{\sigma}$ (defined in 2.3), the connected component of the dependency graph containing $v$ after reducing $\Phi$ under partial assignment $\sigma$. The existence of a fast sampling algorithm, $\mathsf{UniformSample}(\cdot)$, for a uniformly random satisfying assignment to $\Phi_{v}^{\sigma}$ follows from the following theorem, which we specialize to our setting.
Theorem 4.3 ([16, Theorems 5.1 and 6.1]).
Suppose $\Phi$ is a $(k,d)$-CNF formula on $n$ variables with $k\cdot 2^{-k}\cdot(dk)^{5}\cdot 4\leq\frac{1}{150e^{3}}$. Then there exists an algorithm $\mathsf{UniformSample}$ that terminates in $O(k^{3}(dk)^{9}n)$ time in expectation, and outputs a uniformly random satisfying assignment for $\Phi$.
Correctness for Algorithm 1 follows from the above correctness of $\mathsf{MarginSample}$ (Proposition 4.2) and $\mathsf{UniformSample}$ (Theorem 4.3) algorithms, along with the independence of marginal distributions in different connected components computed by $\mathsf{Conn}$.
4.2 Efficiency
In this section, we discuss the efficiency of Algorithm 1. We start with the following proposition that gives an upper bound on the size of $\mathsf{Conn}(\Phi,\sigma,v)$ in Algorithm 3, where $v$ is a marked variable.
Proposition 4.4.
Suppose Algorithm 3 is called during the execution of $\mathsf{MarginSample}(\Phi,\emptyset,v)$ with $\mathsf{IsMarked}(v)=1$. Let $\Phi^{\sigma}_{v}=\mathsf{Conn}(\Phi,\sigma,v)$ in line 9 of Algorithm 3. Suppose $d\leq 2^{\alpha k/4}$. Then with probability at least $1-n^{-0.1\log n}$ the number of clauses in $\Phi^{\sigma}_{v}$ is at most $kd\log^{2}n$.
Proof.
Let $\Psi=(V_{\Psi},\mathcal{C}_{\Psi})$ denote the induced $k$-CNF obtained by taking all the clauses in $\Phi$ in their entirety that appear (possibly in a reduced form) in $\Phi^{\sigma}_{v}$. In particular $\Psi$ is $k$-uniform (although $\Phi^{\sigma}_{v}$ might not be), and $\Phi_{v}^{\sigma}$ and $\Psi$ have the same number of clauses. Let $\mathcal{M}_{\Psi}$ be the set of marked variables in $V_{\Psi}$.
For any $T\subseteq\mathcal{M}_{\Psi}$, we say that $T$ is a 2-tree if both of the following conditions are satisfied:
  1. 1.
    for all $u,v\in T$, $\text{dist}_{H_{\Phi}}(u,v)\geq 2$ (where $H_{\Phi}$ is the dependency hypergraph of 2.1);
  2. 2.
    if one adds an edge between every $u,v$ with $\text{dist}_{H_{\Phi}}(u,v)=2$, then $T$ is connected.
Suppose $\Psi$ has $b$ clauses. Then by a greedy removal argument, one can find a 2-tree $T\subseteq\mathcal{M}_{\Psi}$ with $|T|=\left\lfloor\frac{b}{kd}\right\rfloor$. For every $w\in T$, let $C_{w}\in C_{\Psi}$ be any clause that contains $w$. We know that clauses in $\{C_{w}:w\in T\}$ have pairwise disjoint variable sets as $\text{dist}_{H_{\Phi}}(u,v)\geq 2$ for all $u,v\in T$.
Since $\Phi_{v}^{\sigma}$ is a single connected component, we know that for every marked variable $w^{\prime}\in\text{vbl}(C_{w})$, $\sigma(w^{\prime})$ does not satisfy $C_{w}$. This occurs with probability $1-\theta$, as we might set $w^{\prime}\leftarrow\star$ which occurs with probability $1-2\theta$, or we might set $w^{\prime}\leftarrow x\in\{0,1\}$ such that $(C_{w})_{w^{\prime}\leftarrow x}\neq 1$, which occurs with probability $\theta$.
Recall that every $C_{w}$ has at least $\alpha k$ marked variables. Moreover, all of these variables have been accessed by the partial assignment $\sigma$, as we call $\mathsf{MarginOverflow}$ on marked variables in a depth first search fashion. Thus, we get that
$\mathbb{P}[\text{$C_{w}$ is not satisfied by $(\sigma(w^{\prime}))_{w^{\prime}\in\text{vbl}(C_{w})\cap\mathcal{M}}$}]\leq(1-\theta)^{\alpha k}.$
Since the $C_{w}$ have disjoint variable sets, the events whether $C_{w}$ are satisfied by $(\sigma(w^{\prime}))_{w^{\prime}\in\text{vbl}(C_{w})\cap\mathcal{M}}$ are independent of each other. Thus, we know that
$\mathbb{P}[\text{none of the $C_{w}$, $w\in T$ is satisfied by $\sigma$}]\leq(1-\theta)^{\alpha k|T|}=(1-\theta)^{\alpha k\left\lfloor\frac{b}{kd}\right\rfloor}.$
Finally, we union bound over all 2-trees of size $\left\lfloor\frac{b}{kd}\right\rfloor$ that contains $v$. It is well-known (see [8, Corollary 5.7]) that, since $H_{\Phi}$ is a $k$-uniform hypergraph with maximum degree $d$, the number of 2-trees $T\subseteq H_{\Phi}$ containing $v$ with $|T|=\ell$ is at most $\frac{(ed^{2})^{\ell-1}}{2}$. Setting $\ell=\left\lfloor\frac{b}{kd}\right\rfloor$ and applying a union bound, we get that
$\mathbb{P}[\text{$\Phi_{v}^{\sigma}$ has $b$ clauses}]=\mathbb{P}[\text{$\Psi$has $b$ clauses}]\leq\frac{(ed^{2})^{\left\lfloor\frac{b}{kd}\right\rfloor-1}}{2}(1-\theta)^{\alpha k\left\lfloor\frac{b}{kd}\right\rfloor}\leq 0.9^{\left\lfloor\frac{b}{kd}\right\rfloor}.$
Here in the last inequality, we used the fact that $\theta>0.4$ implied by 3 and the assumption that $d\leq 2^{\alpha k/4}$. Setting $b=kd\log^{2}n$, we get that $\mathbb{P}[\text{$\Phi_{v}^{\sigma}$ has $b$ clauses}]\leq 0.9^{\log^{2}n}\leq n^{-0.1\log n}.$
With Proposition 4.4, we show that Algorithm 1 takes $\mathrm{polylog}(n)$ time in expectation for every marked variable.
Proposition 4.5.
Suppose $d\leq 2^{\alpha k/4}$. Let $\Phi$ be a $(k,d)$-formula satisfying Lemma 2.11. Suppose $(\mathsf{IsMarked}(v))_{v\in V}$ gives a valid marking $\mathcal{M}$. Then for all $v\in\mathcal{M}$, Algorithm 1 takes $\mathrm{polylog}(n)$ time in expectation to compute.
Proof.
Suppose we run Algorithm 1 on a marked variable $v$. Then we have the following breakdown of the complexity of Algorithm 1:
  1. 1.
    Computing $\mathsf{IsMarked}(v)$. By Theorem 2.12, this takes $\mathrm{polylog}(n)$ time in expectation.
  2. 2.
    In $\text{MarginSample}(\Phi,\emptyset,v)$, we first sample $r_{v}\in[0,1]$ uniformly at random.
    • If $r_{v}\leq 2\theta$, then we choose $\sigma(v)\in\{0,1\}$ uniformly at random, which takes constant time.
    • If $r_{v}>2\theta$, then we call $\mathsf{MarginOverflow}(\Phi,\emptyset_{v\leftarrow\star},v)$. Let $s$ denote the number of clauses in $\Phi_{v}^{\sigma}=\mathsf{Conn}(\Phi,\sigma,v)$ in line 9 of $\mathsf{MarginOverflow}(\Phi,\emptyset_{v\leftarrow\star},v)$. By Proposition 4.4, we know that
      $\mathbb{E}[s]\leq n^{-0.1\log n}|\mathcal{C}|+(1-n^{-0.1\log n})kd\log^{2}n\leq n^{-0.1\log n}dn+kd\log^{2}n\leq 2kd\log^{2}n.$
      Given that $\Phi_{v}^{\sigma}$ has $s$ clauses, we know that there are at most $O(s)$ further iterations of $\mathsf{MarginOverflow}$ within $\mathsf{MarginOverflow}(\Phi,\emptyset_{v\leftarrow\star},v)$ as $\Phi$ has constant maximum degree. Each iteration of $\mathsf{MarginOverflow}$ involves computing $\mathsf{IsMarked}(w)$ for a constant number of variables $w$ (plus a constant number of random samplings), so by Theorem 2.12, it takes $\mathrm{polylog}(n)$ time in expectation. Summing up, we know that computing $\Phi_{v}^{\sigma}$ takes $\mathrm{polylog}(n)\cdot s$ time in expectation. Finally, by Theorem 4.3, we know that it takes $O(s)$ time in expectation to compute $\mathsf{UniformSample}(\Phi_{v}^{\sigma})$. This altogether tells us that $\mathsf{MarginOverflow}(\Phi,\emptyset_{v\leftarrow\star},v)$ has expected complexity at most $\mathrm{polylog}(n)$.
We now show that Algorithm 1 takes $\mathrm{polylog}(n)$ time in expectation for every unmarked variable as well.
Proposition 4.6.
Suppose $d\leq 2^{\alpha k/4}$. Let $\Phi$ be a $(k,d)$-formula satisfying Lemma 2.11. Suppose $(\mathsf{IsMarked}(\Phi,v))_{v\in V}$ gives a valid marking $\mathcal{M}$. Then, for all $v\in V\setminus\mathcal{M}$, Algorithm 1 takes $\mathrm{polylog}(n)$ time in expectation to compute.
Proof.
Let $\sigma\in\{0,1\}^{\mathcal{M}}$ denote the partial assignment Algorithm 1 would output on $\mathcal{M}$, were we to compute $\sigma_{v}$ for every $v\in\mathcal{M}$. By Theorem 4.1, we know that $\sigma$ follows the marginal distribution $\mu_{\mathcal{M}}$. Moreover, we know from Lemma 2.11 that with probability at least $1-\delta/n^{2}$, $\sigma$ $\delta$-shatters $\Phi$, which means that every connected component in $\Phi^{\sigma}$ has size $\lesssim dk\log(n/\delta)$. We will not actually construct the entire partial assignment $\sigma$, but will locally use this shattering property as described below.
For every $v\in V\setminus\mathcal{M}$, let $\Phi^{\sigma}_{v}$ denote the connected component in $\Phi^{\sigma}$ containing $v$. Note that $\Phi^{\sigma}_{v}=\mathsf{Conn}(\Phi,\emptyset,v)$ and thus can be locally computed without extending $\sigma$ to a full assignment.
Suppose $\Phi^{\sigma}_{v}$ has $s$ variables. By Theorem 2.12 and Proposition 4.5, both $\mathsf{IsMarked}(\cdot)$ and $\mathsf{MarginSample}(\Phi,\emptyset,w)$ (for any $w\in\mathcal{M}$) have expected $\mathrm{polylog}(n)$ complexity. Since we execute each of these two operations $O(s)$ times when we compute $\Phi_{v}=\mathsf{Conn}(\Phi,\emptyset,v)$, we know that $\mathsf{Conn}(\Phi,\emptyset,v)$ has expected $s\cdot\mathrm{polylog}(n)$ complexity. It remains to sample a uniformly random solution using $\mathsf{UniformSample}(\Phi^{\sigma}_{v})$. By Theorem 4.3, we know that $\mathsf{UniformSample}(\Phi^{\sigma}_{v})$ has expected $O(s)$ complexity.
Since $s\lesssim dk\log^{2}n$ with probability at least $1-\delta/n^{2}$, we get that in expectation,
  • $\mathsf{Conn}(\Phi,\sigma,v)$ has time complexity at most
    $(1-\delta/n^{2})\cdot dk\log^{2}n\cdot\mathrm{polylog}(n)+\delta/n^{2}\cdot n\cdot\mathrm{polylog}(n)=\mathrm{polylog}(n);$
  • $\mathsf{UniformSample}(\Phi^{\sigma}_{v})$ has complexity at most
    $(1-\delta/n^{2})\cdot O(dk\log^{2}n)+\delta/n^{2}\cdot O(n)=\mathrm{polylog}(n).$
Therefore, Algorithm 1 takes $\mathrm{polylog}(n)$ time in expectation to compute. ∎
4.3 Completing the proof
Having gathered all the ingredients, we can now prove the main theorem that is the formal version of 1.1.
Theorem 4.7.
Suppose $\Phi=(V,\mathcal{C})$ is a $(k,d)$-formula satisfying (1). Let $\mu$ be the uniform distribution over satisfying assignments to $\Phi$, with marginal distributions $\mu_{v}$ for $v\in V$. Then for all $c>0$, there is a $(\mathrm{polylog}n,n^{-c})$-random local access algorithm $\mathcal{A}$ for sampling the variable assignment of $v\in V$ as $\widehat{\mu}_{v}$, such that
$d_{\operatorname{\mathsf{TV}}}(\mathsf{Law}(\widehat{{\mu}_{v}}),\mu_{v})\leq\frac{1}{n^{c}}.$
Proof.
Suppose $\Phi=(V,\mathcal{C})$ is a $(k,d)$-formula satisfying (1). Fix constant $c>0$.
We first choose the parameters $\alpha,\beta_{1},\beta_{2}$ that satisfy Theorem 2.12. Let
$\alpha=1/75,\qquad\beta_{1}=0.778,\qquad\beta_{2}=0.96.$
One can verify that $1/2<\beta_{1}<\beta_{2}<1-\alpha$ and with these parameters, conditions in Theorem 2.12 are satisfied. Thus, by Theorem 2.12, there exists a $\mathrm{polylog}(n)$ oblivious local computation algorithm $\mathsf{IsMarked}(\cdot)$ that with probability at least $1-1/n$ gives a consistent $\alpha$-marking $\mathcal{M}\subseteq V$.
Suppose $\mathsf{IsMarked}(\cdot)$ gives a consistent $\alpha$-marking $\mathcal{M}\subseteq V$. Since $d\lesssim 2^{k/400}$, condition 4 is satisfied by parameters $\alpha,d,k$. By Proposition 4.2 and Theorem 4.3, we know that for all variable $v\in V$, Algorithm 1 on $v$ returns some $\widehat{\mu}_{v}$ within total variation distance $n^{-c}$ of $\mu_{v}$.
Since $d\lesssim 2^{k/400}$, our choice of $\alpha$ gives the condition $d\leq 2^{\alpha k/4}$ required by Propositions 4.5 and 4.6. These two propositions tell that Algorithm 1 on every $v\in V$ takes $\mathrm{polylog}(n)$ time in expectation.
By construction, Algorithm 1 is memory-less as it samples all necessary variable assignments in order to compute the assignment of a queried variable $v$. Furthermore, Algorithm 1 queried on different variables $v\in V$ collectively returns an assignment $\sigma\in\Omega$ that has $d_{\operatorname{\mathsf{TV}}}(\textsf{Law}(\sigma),\mu)<n^{-(c-1)}$ by construction. Since this holds for any $c>0$ constant, we obtain the desired result. ∎
5 Concluding remarks
With more involved refinements and optimizations of the arguments in this work, the density constraint $d\lesssim 2^{k/400}$ of 1.1 can be substantially improved (to something closer to $d\lesssim 2^{k/50}$). We omit these additional calculations in favor of expositional clarity to highlight our main result: random local access models exist for arbitrary bounded degree $k$-CNFs at exponential clause density. Furthermore, these arguments can also be adapted (in a similar fashion to [11, 17, 5]) to obtain a random local access models for random $k$-CNFs in a comparable density regime.
Nonetheless, the limit of the approaches in this work would still fall well short of obtaining random local access for, e.g., approximately $d\lesssim 2^{k/5}$, the largest density at which we currently know how to efficiently sample solutions to an arbitrary bounded degree $k$-CNF in nearly-linear time ([16]). This is because of our reliance on a query-oblivious LCA to construct a local marking and our application of sampling results to a correspondingly reduced CNF.
The approach we take in this work is only one of many possible schema to reduce from existing classical, parallel, and/or distributed algorithms to more local algorithms. Our approach involved using ideas and techniques from a variety of previous works (including [1, 16, 21]), many of which were in the non-local setting, and adapting and combining them in a novel way to obtain a sublinear sampler. In [4], the authors adapted a parallel Glauber dynamics algorithm for sampling colorings to obtain random local access to random proper $q$-colorings. We conjecture that one could, using similar ideas to [4], adapt parallel Glauber dynamics or algorithmic correlation decay (as studied in [9, 20]) in conjunction with our local marking scheme to obtain different sampling algorithms. We expect such algorithms would also provide random local access to satisfying assignments of a $k$-CNF at comparable clause density to this work.
Appendix A Proof of Theorem 2.12
Given a $k$-uniform hypergraph $H$ and constants $10<\alpha<\beta_{1}<\beta_{2}<$ 1, we describe an algorithm $\mathcal{A}$ that finds a 2-coloring of $H$, such that every edge in $H$ has at least $\alpha k$ vertices in each color. Taking $H=H_{\Phi}$ (as in 2.1), we see that finding an $\alpha$-marking for $\Phi$ is equivalent to finding such a 2-coloring for $H_{\Phi}$. The algorithm $\mathcal{A}$ runs in three phases.
  • In Phase 1, we randomly color each vertex following an arbitrary order of the vertices.
    • If an edge at any point contains at least $\alpha k$ vertices in both colors, then we delete it from $H$.
    • If an edge at any point contains more than $\beta_{1}k$ vertices in one color, then we call it a dangerous-1 edge and mark all its uncolored vertices as troubled-1. These troubled-1 vertices will not be colored in Phase 1.
    • If an edge at any point has all its vertices colored or troubled-1, but the edge itself is not deleted or dangerous-1, then we call this edge unsafe-1.
    We proceed until all vertices in $H$ become colored or troubled-1.
  • In Phase 2, we run the previous phase again on the remaining hypergraph $H$. There will be two differences:
    • We do not recolor vertices that are already colored in Phase 1. In other words, we only randomly color those troubled-1 vertices.
    • This time, we call an edge dangerous-2 if it contains more than $\beta_{2}k$ vertices in one color. Again, if an edge becomes dangerous-2 at any point, then we mark all its uncolored vertices as troubled-2. These troubled-2 vertices will not be colored in Phase 2.
    We proceed until all vertices in $H$ become colored or troubled-2.
  • In Phase 3, we use an exhaustive search to color all the remaining (i.e., troubled-2) vertices in $H$, such that every remaining edge in $H$ has at least $\alpha k$ vertices in each color.
Suppose $1/2<\beta_{1}<\beta_{2}<1-\alpha$ satisfy conditions Theorem 2.12:
$4\alpha<2(1-\beta_{2})<1-\beta_{1},$
$16k^{4}d^{5}\leq 2^{(\beta_{1}-h(1-\beta_{1}))k},$
$16k^{4}d^{5}\leq 2^{(\beta_{2}-\beta_{1})k-h(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}})(1-\beta_{1})k},$
$\delta\mapsto(\beta_{2}-\delta)-h(\frac{\beta_{2}-\delta}{1-\delta})(1-\delta)\text{ is decreasing on $0\leq\delta\leq\beta_{1}$,}$
$2e(kd+1)<2^{(1-h(\frac{\alpha}{1-\beta_{2}}))(1-\beta_{2})k}.$
Then, the following holds regarding the above algorithm.
Lemma A.1.
With probability at least $1-\frac{1}{n}$, after Phase 1, all connected components in $H$ have size at most $c_{1}\log n$.
Proof.
Let $H^{3}$ deote the distance-3 graph of $H$; that is, we have $V(H^{3})=V(H)$, and $u\sim v$ in $H^{3}$ if they have distance 3 in $H$. We say that $W\subseteq V(H)$ is a 3-tree if $H^{3}[W]$ is connected.
Observe that for every $w>0$, there are at most $n4^{w}(dk)^{3w}$ 3-trees on $w$ vertices. This is because $H^{3}$ has maximum degree at most $(dk)^{3}$. Since there are at most $4^{w}$ non-isomorphic trees on $w$ vertices, and at most $n(dk)^{3w}$ copies of every such tree in $H^{3}$, we get that there are at most $n4^{w}(dk)^{3w}$ 3-trees on $w$ vertices.
Let $e$ be any edge, and $v$ be any vertex in $H$. Observe that
[…]
$\leq 4kd^{2}\binom{k}{\beta_{1}k}2^{-\beta_{1}k}\leq 4kd^{2}2^{(h(1-\beta_{1})-\beta_{1})k}.$
Let $W$ be a 3-tree on $w$ vertices, and $p=4kd^{2}2^{(h(1-\beta_{1})-\beta_{1})k}$. Then
$\mathbb{P}[\text{$W$ is contained in some connected component after Phase 1}]\leq p^{w}.$
Thus, the expected number of 3-trees on $w$ vertices after Phase 1 is at most
$n4^{w}(dk)^{3w}p^{w}=n(4d^{3}k^{3}\cdot 4kd^{2}2^{(h(1-\beta_{1})-\beta_{1})k})^{w}=n(16k^{4}d^{5}\cdot 2^{(h(1-\beta_{1})-\beta_{1})k})^{w},$
which is at most $1/n$ when $w=c_{1}\log n$. By Markov’s inequality, with probability at least $1-1/n$, there is no 3-tree of size larger than $c_{1}\log n$. Since every size-$s$ connected component in $H$ contains a 3-tree of size at least $s/(dk)^{3}$, we get that with probability at least $1-1/n$, there is no connected component of size larger than than $c_{1}(dk)^{3}\log n$ in $H$. ∎
Lemma A.2.
Let $H_{1}$ be a connected component of $H$ after Phase 1. Suppose $H_{1}$ has size at most $c_{1}\log n$. Then with probability at least $1-\frac{1}{\log^{2}n}$, after Phase 2, all connected components in $H_{1}$ have size at most $c_{2}\log\log n$.
Proof.
Let $\widetilde{H}_{1}$ be the hypergraph obtained by deleting all colored vertices from $H_{1}$. Suppose $N\leq c_{1}\log n$ is the vertex number of $\widetilde{H}_{1}$. Note that $\widetilde{H}_{1}$ also has maximum degree at most $d$, and $\widetilde{H}_{1}^{3}$ also has maximum degree at most $(dk)^{3}$.
The argument in similar to the proof of Lemma A.1. Let $e$ be any edge, and $v$ be any vertex in $\widetilde{H}_{1}$. Observe that
$\mathbb{P}[\text{$e$ becomes dangerous-2 in Phase 2}]\leq 2\binom{(1-\delta)k}{(\beta_{2}-\delta)k}2^{-(\beta_{2}-\delta)k}\text{ for some $0\leq\delta\leq\beta_{1}$},$
$\mathbb{P}[\text{$v$ becomes troubled-2 in Phase 2}]\leq 2d\binom{(1-\delta)k}{(\beta_{2}-\delta)k}2^{-(\beta_{2}-\delta)k},$
$\mathbb{P}[\text{$e$ becomes unsafe-2 in Phase 2}]\leq 2kd\binom{(1-\delta)k}{(\beta_{2}-\delta)k}2^{-(\beta_{2}-\delta)k},$
[…]
$\qquad\leq 4kd^{2}\binom{(1-\delta)k}{(\beta_{2}-\delta)k}2^{-(\beta_{2}-\delta)k}$
$\qquad\leq 4kd^{2}2^{h(\frac{\beta_{2}-\delta}{1-\delta})(1-\delta)k-(\beta_{2}-\delta)k}\leq 4kd^{2}2^{h(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}})(1-\beta_{1})k-(\beta_{2}-\beta_{1})k}.$
Let $\widetilde{p}=4kd^{2}2^{h(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}})(1-\beta_{1})k-(\beta_{2}-\beta_{1})k}$. Thus, the expected number of 3-trees on $w$ vertices after Phase 1 is at most
$N4^{w}(dk)^{3w}\widetilde{p}^{w}$ $=N(4d^{3}k^{3}\cdot 4kd^{2}2^{h(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}})(1-\beta_{1})k-(\beta_{2}-\beta_{1})k})^{w}$
$=N(16k^{4}d^{5}\cdot 2^{h(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}})(1-\beta_{1})k-(\beta_{2}-\beta_{1})k})^{w},$
which is at most $1/N^{2}$ when $w=c_{2}\log N$. By Markov’s inequality, with probability at least $1-1/N^{2}$, there is no 3-tree of size larger than $c_{2}\log N$. This implies that with probability at least $1-1/\log^{2}n$, there is no connected component of size larger than than $c_{2}(dk)^{3}\log\log n$ in $H$. ∎
Lemma A.3.
Let $H_{2}$ be a connected component of $H$ after Phase 2. Suppose $H_{2}$ has size at most $c_{2}\log\log n$. Then there exists a valid coloring of $H_{2}$, i.e., a coloring in which every edge has at least $\alpha n$ vertices under each color.
Proof.
Consider a random 2-vertex coloring of $H_{2}$. For every edge $e$ in $H_{2}$, let $A_{e}$ be the event that $e$ does not have $\alpha n$ vertices under one of the colors. Then we have
$\mathbb{P}[A_{e}]\leq 2\binom{(1-\beta_{2})k}{\leq\alpha k}2^{-(1-\beta_{2})k}\leq 2^{(h(\frac{\alpha}{1-\beta_{2}})-1)(1-\beta_{2})k+1}.$
Since each $A_{e}$ is independent from all but at most $kd$ other $A_{e^{\prime}}$’s, with
$e\cdot 2^{(h(\frac{\alpha}{1-\beta_{2}})-1)(1-\beta_{2})k+1}\cdot(kd+1)<1,$
we know from the Lovász local lemma that with positive probability, none of the $A_{e}$’s occurs. ∎
We now argue that for every $v\in V(H)$, we can locally access the color of $v$ given by the above algorithm. To do so, observe that the above algorithm works for an arbitrary variable ordering. In particular, a random ordering will work (which we can locally look up given a public tape of random bits $\mathbf{R}$).
We associate a random value $a_{v}\in[0,1]$ to every vertex $v\in V(H)$, and order the vertices as $v_{1},\dots,v_{n}$ such that $a_{v_{1}}<\dots<a_{v_{n}}$. For every vertex $v\in V$, we compute the query tree $\mathcal{T}_{v}$ defined as follows:
  • Let $v$ be the root of $\mathcal{T}_{v}$.
  • For every $w\in V$ that shares a clause with $v$, if $a_{w}<a_{x}$, then we add $w$ to $\mathcal{T}_{v}$ as a child of $v$.
  • We build $\mathcal{T}_{v}$ in this way recursively until no child node can be added to $\mathcal{T}_{v}$.
We know that with high probability, $\mathcal{T}_{v}$ has $\mathrm{polylog}(n)$ vertices.
Lemma A.4 ([1, Theorem 3.1]).
For every $k,d\geq 1$, there exists $c_{3}>0$ such that for all sufficiently large $n$, $\mathbb{P}[|\mathcal{T}_{v}|>c_{3}\log^{kd+1}(n)]<1/n^{2}$.
In Phase 1, whether $v$ is colored or troubled-1 completely depends on its descendants in $\mathcal{T}_{v}$. Thus, to compute the assignment of $v$ in Phase 1, it suffices to run Phase 1 on all vertices in $\mathcal{T}_{v}$. By Lemma A.4, with probability $1-1/n^{2}$, this step takes $\mathrm{polylog}(n)$ time.
If $v$ gets colored in Phase 1, then we return the color directly. Otherwise, if $v$ becomes troubled-1 after Phase 1, then we compute the connected component of $H$ containing $v$ after Phase 1. Recall from Lemma A.1 that with probability $1-1/n$, all connected components after Phase 1 has size at most $c_{1}\log n$, so this step takes another $\mathrm{polylog}(n)$ time.
Let $H_{v}$ denote this connected component. By Lemma A.2, if we perform $O(\log n/\log\log n)$ parallel instances of Phase 2 on $H_{v}$, then with probability $1-1/n^{2}$, in one of them $H_{v}$ is further reduced to connected components each of size $c_{2}\log\log n$. We then perform Phase 3 to find a valid coloring on the component that contains $v$, which by Lemma A.3 is guaranteed to succeed and finishes in $2^{O(\log\log n)}=\mathrm{polylog}(n)$ time.
Combining the above analysis, we have obtained the local computation algorithm $\mathsf{IsMarked}(\cdot)$ with the desired properties.